# LLM-Craft: Robotic Crafting of Elasto-Plastic Objects with Large Language Models

### v1

@@ Abstract
When humans create sculptures, we are able to reason about how geometrically we need to alter the clay state to reach our target goal. We are not computing point-wise similarity metrics, or reasoning about low-level positioning of our tools, but instead determining the higher-level changes that need to be made. In this work, we propose LLM-Craft, a novel pipeline that leverages large language models (LLMs) to iteratively reason about and generate deformation-based crafting action sequences. We simplify and couple the state and action representations to further encourage shape-based reasoning. To the best of our knowledge, LLM-Craft is the first system successfully leveraging LLMs for complex deformable object interactions. Through our experiments, we demonstrate that with the LLM-Craft framework, LLMs are able to successfully reason about the deformation behavior of elasto-plastic objects. Furthermore, we find that LLM-Craft is able to successfully create a set of simple letter shapes. Finally, we explore extending the framework to reaching more ambiguous semantic goals, such as ”thinner” or ”bumpy”. For videos please see our website: https://sites.google.com/andrew.cmu.edu/llmcraft.
Keywords: Deformable Object Manipulation, Large Language Models, Long Horizon Planning
1 Introduction
@@
With the goal of building robust and generalizable autonomous robotic systems, we need to design frameworks that can reason about complex interactions between the robot and the environment. Deformable object manipulation is a great test task for reasoning about these complex interactions, as it requires understanding how the state of the object itself will change during contact. In this work, we will investigate the task of deformation-based crafting, in which the robot moulds the clay into a variety of simple shapes with a parallel gripper. This task allows us to investigate the challenges of both predicting the behavior of the clay as well as long-horizon planning and reasoning. Previous work focused on elasto-plastic object manipulation and crafting typically work at the very low-level, predicting clay dynamics [1, 2, 3], or direct action imitation [4]. However, these approaches do not leverage existing world knowledge to reason about how clay behaves, or have the ability to reason at a geometric or semantic level.
@@
In this work, we present LLM-Craft, a preliminary study exploring how large language models (LLMs) reason about the complex task of elasto-plastic object crafting. LLMs have been demonstrated to contain useful world knowledge for a wide range of reasoning and generation tasks [5, 6, 7, 8, 9, 10]. We hypothesize that LLMs have substantial and useful world knowledge for the crafting task, such as understanding how clay will behave when pressed, and understanding semantic shapes and goals. To better facilitate higher-level geometric reasoning for crafting action selection, we simplify our state and action representation substantially.
Through extensive real-world experiments, we demonstrate the success of leveraging LLMs directly for the clay crafting task. We find that the ability of LLMs to reason at the geometric and semantic level is particularly powerful and presents an argument for future crafting pipelines incorporating LLMs. The key contributions of this work are as follows.
@@ 2 Related Work
Elasto-Plastic Object Manipulation: There have been a variety of successful past works within the realm of elasto-plastic object manipulation and crafting. There are numerous simulations modeling the deformation behavior of these objects, but there remains a large sim-to-real gap [11, 12, 13, 14]. These simulations enabled trajectory optimization-based approaches that were able to bridge the sim-to-real gap for simple tasks [15, 16]. Many real-world clay crafting frameworks learn the low-level dynamics of the clay to be used to generate action trajectories [1, 3, 2, 17, 18]. Alternatively, a range of imitation-based frameworks have been successful, where demonstrations are abstracted into skills [19, 20], or more direct trajectory imitation [4]. All of these successful real-world approaches require hardware and task-specific datasets, with many requiring human demonstrations which are particularly time consuming to collect. Additionally, these methods are very brittle to particular design choices and are not always transferable across lab setups. In this work, we investigate if we can leverage the general world knowledge of LLMs to shape elasto-plastic objects with a parallel gripper. From the lens of the deformable manipulation community, if LLMs do indeed have useful world knowledge to make high-level predictions of the objects’ behavior, this could be a useful component to future pipelines that could improve generalizability.
@@
LLMs for Robotic Tasks: There is a significant body of previous work leveraging LLMs for a variety of robotic tasks. LLMs have been particularly effective as robotic task planning agents [21, 22, 23, 24, 25, 26, 27, 28, 29, 30] with many recent works finding success incorporating a variety of real-world grounding techniques including scene graphs [24], diffusion models [26], counterfactual perturbation [31], developing a multimodal language model for embodied tasks [32] and 3D value maps [33]. These LLM-based frameworks span a wide range of reasoning levels, from low-level autoregressive code policy generation [34] to reasoning about task plans purely with natural language [35]. While the widespread success of LLM robotic methods has been impressive, most research has focused on the relatively simple tasks of navigation [25, 36, 24, 37], or sequences of pick and place actions [27, 26, 21, 23, 29]. These tasks do require long-horizon reasoning as well as recognition and interaction with a variety of real-world objects, but the general action skillset is relatively simple and does not require much intuition and understanding of physical robot-object interactions. In Tidybot [38], researchers present an LLM planning framework that involves interacting with deformable cloth, but the specific sorting tasks explored do not require reasoning about the complex deformable dynamics and the robot-object interactions remain only pick and place. Similarly, in [39] researchers explore how LLMs can creatively reason about tool use to achieve a set of tasks, but the tool-object interactions remain very simple. In fact, in [40] researchers concluded that LLMs often simplify or incorrectly predict more complicated interactions between the robot and objects. However, the general motivation behind the significant body of LLM-based robotic methods is that they are leveraging the world knowledge that these LLMs contain to reason about the task. In this work, we hypothesize and demonstrate that LLMs do indeed have the ability to predict and reason about complex robot-object interactions with the task of the robotic crafting. We particularly explore how LLMs can plan complicated action sequences that require reasoning about the effect of the grasp action on the objects’ state itself.
3 Method
@@
The key components of LLM-Craft are the choice of state and action representations as well as the iterative prompting scheme with the LLM. An overview of the pipeline is shown in Figure 1. In this work, we are using Gemini 1.0 Pro Vision [41] as our multimodal large language model that can take both images and text as input. We chose this version of Gemini over the more recent Gemini 1.5 [42] models as Gemini 1.0 Pro Vision was optimized for generating image descriptions and identifying objects in images which is a key part of the reasoning task for crafting shapes. Additionally with preliminary experiments we found it to substantially outperform Gemini 1.5. We have not investigated alternative LLMs, but the LLM-Craft pipeline can be used with any LLM that can take both images and text as input, such as GPT-4 [43].
3.1 State and Action Representation
@@
In this work we are proposing an LLM-based framework for the task of top-down shape crafting of elasto-plastic objects with a parallel gripper, meaning the only actions the system is able to apply to the clay is a squeeze of a fixed distance between the fingertips. In existing robotic clay crafting literature, the common choice of state representation is a point cloud, and the action representation is the continuous end-effector pose [1, 2, 3, 4]. While these state and action choices are very effective in existing pipelines, they can disconnect the reasoning about actions from the state representation. When reasoning about how to shape clay, humans determine at a high level where and how the clay needs to be deformed. In this work, we propose a simple gridded top-down image of the clay as the state and action space to allow the LLM to reason about the larger geometric regions of the clay without needing to directly reason about global end effector poses.
@@
3.2 Prompting
The goal of this work is to explore how we can leverage the world knowledge and shape-level reasoning capabilities of LLMs without requiring any further training or fine-tuning on our specific task. The full process diagram of the iterative prompting scheme is shown in Figure 1, in which we use an action prompt and a termination prompt. The role of the action prompt is to query the LLM with an explanation of the task, state and action space, and the robot embodiment to then output a crafting action sequence of varying length. We take inspiration from past successful approaches using LLMs for robotic tasks for some of the prompting details [40]. The components of the action prompt are as follows: (1) an overview explaining the robot embodiment and the high-level details of the crafting task, (2) an explanation of the environmental setup including the clay and the grid coordinate system, (3) an explanation of the grasp action, including that the end-effector pose cannot change during the grasp, (4) an explanation of how to generate a crafting trajectory, including the requirements of (a) describing the current and goal shapes, (b) reasoning about the grasp strategy, (c) describe the predicted effect of the chosen grasp to the clay state, (5) the goal gridded image, (6) the current state gridded image, and (7) the user command to create the goal from the current state. The role of the termination prompt is to query the LLM to determine if the goal shape has been reached. The advantage of using an LLM as opposed to similarity metric thresholding is that we are prompting the LLM to judge the shape similarity more holistically, i.e. if both of the shapes are of the letter T, which is a fundamentally different task from judging point-wise similarity between point clouds. The terimation prompt consists of (1) an overview of the shape comparison task, (2) an explanation of how to generate a decision, with the requirements to (a) describe the similarities and differences between the state and goal, and (b) make a decision if crafting should continue, (3) the current gridded state image, (4) the goal gridded image, and (5) the user command to determine if crafting should stop. For full details of both prompts, please see the Supplemental Materials. In section 4.4 we will discuss prompt ablations to evaluate the impactfullness of each component of the prompts on the final system performance.
4 Experiments
@@
4.1 State Representation
4.2 Single-Step Goals
@@
Before we leverage the LLM for longer-horizon shape crafting goals, we first need to evaluate how well an LLM can out-of-the-box select the single action to achieve a single-step goal. While this is a relatively simple task, it proved somewhat difficult for the LLM due to how it is reasoning about the clay moulding task. We found that the LLM prescribes semantically meaningful descriptions for the clay states and goals in order to generate action sequences. This is a very useful attribute for the long horizon shape crafting task, but can interfere with performance for the single-step goals in which the goal shape isn’t anything in particular. For more detailed analysis of the LLM reasoning, please see the Supplemental Materials. We found the LLM was able to successfully reach the goal shape, but could sometimes get confused about rotation. The numerical results for the single-step experiments are shown in Table 1, and the qualitative results are visualized in Figure 3.
4.3 Long-Horizon Shape Goals
Loop | CD | EMD |
N.R. | 0.0056 0.0019 | 0.0082 0.0026 |
Iter. | 0.0045 0.0015 | 0.0067 0.0020 |
@@ 4.3 Long-Horizon Shape Goals
To further investigate the capabilities of LLMs for crafting, we evaluated the full system performance on a set of long-horizon letter shape goals consisting of the shapes C, I, L, T, X. We selected these letters because they are all able to be created with only deformation actions, whereas letters such as P are not because our current action formulation cannot create holes. To determine the effectiveness of the proposed iterative prompting scheme, we compared the performance with replanning enabled compared to executing the entire LLM-generated trajectory without replanning. Additionally, to disambiguate the system errors from the limitations of the coarsely discretized action space, we conducted a set of human baseline experiments in which the human operator must select the two cells to squeeze together. The qualitative results are visualized in Figure 4, and the quantitative results are shown in Table 2. We find that the LLM-Craft framework is able to successfully create the simple letter shaped sculptures at a comparable level to our human baseline with the discrete action space. However, it is important to note that the performance of LLM-Craft has a high variance, as sometimes the system selects a bad action early into the trajectory and even with the iterative re-planning framework, the system is unable to recover.
4.4 Prompt Ablation
@@
To evaluate the impact of different components of the prompts on the system performance, we conducted an ablation study removing different parts. In particular, we chose to remove the following: (1) asking the LLM to predict the effect of the grasp on the clay after selecting an action, (2) asking the LLM to describe the similarities and differences between the state and goal before generating an action sequence, (3) removing the termination prompt altogether, (4) removing the goal text description, i.e. if the goal is an L to not mention L in the prompt explicitly, (5) removing the explanation of grasp types, specifically horizontal, vertical, or diagonal, (6) removing the instructions to choose action step-by-step, and (7) removing the explanations of the gridded state and action space. For the complete prompt and prompt variations for each ablation experiment, please see the Supplemental Materials. The breakdown of earth mover’s distance for the L shape across the ablations is shown in Figure 5. We chose to conduct the ablation experiments with the L shape goal because it is difficult to successfully create the bottom perpendicular line in the L given the discrete action space. Therefore, we believe this goal shape will help illuminate behavior differences given the changes to the prompt. Based on our ablation experiments, it is clear that removing the explanation of the grid from the prompt negatively impacts performance the most. This is an expected result, as the choice of state and action space in LLM-Craft is unique and requires a clear explanation, particularly of the labeling system of the grid in order to select the grid cells for actions. Removing the step-by-step instruction also significantly impacted performance, following the results from [40]. We found that the grasp orientation description, goal text description and the termination prompt all had a mid-sized impact on the final performance of LLM-Craft. Finally, the requirements to reason about the state and goal as well as the predicted effect before selecting a grasp action had minimal impact on the average performance of LLM-Craft, but the variance was substantially reduced with their inclusion in the prompt. This makes sense, as requiring the LLM to reason about the differences between the current state and goal as well as describe physically how a grasp will change the state has the potential to reduce the frequency of physically unreasonable actions to be chosen. However, it is important to highlight that while all these components of the prompt reduce the unreasonable actions selected by the LLM, LLM-Craft does sometimes select actions and the explanation of the predicted effect is not fully physically accurate.
4.5 Semantic Goals

### v2

@@ Abstract
When humans create sculptures, we are able to reason about how geometrically we need to alter the clay state to reach our target goal. We are not computing point-wise similarity metrics, or reasoning about low-level positioning of our tools, but instead determining the higher-level changes that need to be made. In this work, we propose LLM-Craft, a novel pipeline that leverages large language models (LLMs) to iteratively reason about and generate deformation-based crafting action sequences. We simplify and couple the state and action representations to further encourage shape-based reasoning. To the best of our knowledge, LLM-Craft is the first system successfully leveraging LLMs for complex deformable object interactions. Through our experiments, we demonstrate that with the LLM-Craft framework, LLMs are able to successfully reason about the deformation behavior of elasto-plastic objects. Furthermore, we find that LLM-Craft is able to successfully create a set of simple letter shapes. Finally, we explore extending the framework to reaching more ambiguous semantic goals, such as ”thinner” or ”bumpy”. For videos and prompting details, please visit our project website: https://sites.google.com/andrew.cmu.edu/llmcraft/home.
I Introduction
@@
With the goal of building robust and generalizable autonomous robotic systems, we need to design frameworks that can reason about complex interactions between the robot and the environment. Deformable object manipulation is a great test task for reasoning about these complex interactions, as it requires understanding how the state of the object itself will change during contact. In this work, we will investigate the task of deformation-based crafting, in which the robot moulds the clay into a variety of simple shapes with a parallel gripper. This task allows us to investigate the challenges of both predicting the behavior of the clay as well as long-horizon planning and reasoning. Previous work focused on elasto-plastic object manipulation and crafting typically work at the very low-level, predicting clay dynamics [shi2022robocraft, bartsch2023sculptbot, shi2023robocook], or direct action imitation [bartsch2024sculptdiff]. However, these approaches do not leverage existing world knowledge to reason about how clay behaves, or have the ability to reason at a geometric or semantic level.
@@
In this work, we present LLM-Craft, a preliminary study exploring how large language models (LLMs) reason about the complex task of elasto-plastic object crafting. LLMs have been demonstrated to contain useful world knowledge for a wide range of reasoning and generation tasks [mirchandani2023large, wu2023smartplay, todd2023level, yu2023language, jadhav2024large, wang2024can]. We hypothesize that LLMs have substantial and useful world knowledge for the crafting task, such as understanding how clay will behave when pressed, and understanding semantic shapes and goals. To better facilitate higher-level geometric reasoning for crafting action selection, we simplify our state and action representation substantially. Through extensive real-world experiments, we demonstrate the success of leveraging LLMs directly for the clay crafting task. We find that the ability of LLMs to reason at the geometric and semantic level is particularly powerful and presents an argument for future crafting pipelines incorporating LLMs. The key contributions of this work are as follows.
@@ II Related works
Elasto-Plastic Object Manipulation: There have been a variety of successful past works within the realm of elasto-plastic object manipulation and crafting. There are numerous simulations modeling the deformation behavior of these objects, but there remains a large sim-to-real gap [heiden2021disect, hu2018moving, huang2021plasticinelab, gu2023maniskill2]. These simulations enabled trajectory optimization-based approaches that were able to bridge the sim-to-real gap for simple tasks [qi2022learning, yamada2024d]. Many real-world clay crafting frameworks learn the low-level dynamics of the clay to be used to generate action trajectories [shi2022robocraft, shi2023robocook, bartsch2023sculptbot, bauer2024doughnet, li2024deformnet]. Alternatively, a range of imitation-based frameworks have been successful, where demonstrations are abstracted into skills [lin2022planning, li2023dexdeform], or more direct trajectory imitation [bartsch2024sculptdiff]. All of these successful real-world approaches require hardware and task-specific datasets, with many requiring human demonstrations which are particularly time consuming to collect. Additionally, these methods are very brittle to particular design choices and are not always transferable across lab setups. In this work, we investigate if we can leverage the general world knowledge of LLMs to shape elasto-plastic objects with a parallel gripper. From the lens of the deformable manipulation community, if LLMs do indeed have useful world knowledge to make high-level predictions of the objects’ behavior, this could be a useful component to future pipelines that could improve generalizability.
@@
LLMs for Robotic Tasks: There is a significant body of previous work leveraging LLMs for a variety of robotic tasks. LLMs have been particularly effective as robotic task planning agents [wu2024mldt, wang2023describe, bhat2024grounding, rana2023sayplan, dorbala2023can, sharan2024plan, singh2023progprompt, ding2023task, gupta2024action, cheng2024empowering] with many recent works finding success incorporating a variety of real-world grounding techniques including scene graphs [rana2023sayplan], diffusion models [sharan2024plan], counterfactual perturbation [wang2024grounding], developing a multimodal language model for embodied tasks [driess2023palm] and 3D value maps [huang2023voxposer]. These LLM-based frameworks span a wide range of reasoning levels, from low-level autoregressive code policy generation [liang2023code] to reasoning about task plans purely with natural language [mikami2024natural]. While the widespread success of LLM robotic methods has been impressive, most research has focused on the relatively simple tasks of navigation [dorbala2023can, shah2023lm, rana2023sayplan, yang2023llm], or sequences of pick and place actions [singh2023progprompt, sharan2024plan, wu2024mldt, bhat2024grounding, gupta2024action]. These tasks do require long-horizon reasoning as well as recognition and interaction with a variety of real-world objects, but the general action skillset is relatively simple and does not require much intuition and understanding of physical robot-object interactions. In Tidybot [wu2023tidybot], researchers present an LLM planning framework that involves interacting with deformable cloth, but the specific sorting tasks explored do not require reasoning about the complex deformable dynamics and the robot-object interactions remain only pick and place. Similarly, in [xu2023creative] researchers explore how LLMs can creatively reason about tool use to achieve a set of tasks, but the tool-object interactions remain very simple. In fact, in [kwon2023language] researchers concluded that LLMs often simplify or incorrectly predict more complicated interactions between the robot and objects. However, the general motivation behind the significant body of LLM-based robotic methods is that they are leveraging the world knowledge that these LLMs contain to reason about the task. In this work, we hypothesize and demonstrate that LLMs do indeed have the ability to predict and reason about complex robot-object interactions with the task of the robotic crafting. We particularly explore how LLMs can plan complicated action sequences that require reasoning about the effect of the grasp action on the objects’ state itself.
III Methodology
@@
The key components of LLM-Craft are the choice of state and action representations as well as the iterative prompting scheme with the LLM. An overview of the pipeline is shown in Figure 2. In this work, we are using Gemini 1.0 Pro Vision [team2023gemini] as our multimodal large language model that can take both images and text as input. We chose this version of Gemini over the more recent Gemini 1.5 [reid2024gemini] models as Gemini 1.0 Pro Vision was optimized for generating image descriptions and identifying objects in images which is a key part of the reasoning task for crafting shapes. Additionally with preliminary experiments we found it to substantially outperform Gemini 1.5. We have not investigated alternative LLMs, but the LLM-Craft pipeline can be used with any LLM that can take both images and text as input, such as GPT-4 [achiam2023gpt].
III-A State and Action Representation
@@
In this work we are proposing an LLM-based framework for the task of top-down shape crafting of elasto-plastic objects with a parallel gripper, meaning the only actions the system is able to apply to the clay is a squeeze of a fixed distance between the fingertips. In existing robotic clay crafting literature, the common choice of state representation is a point cloud, and the action representation is the continuous end-effector pose [shi2022robocraft, bartsch2023sculptbot, shi2023robocook, bartsch2024sculptdiff]. While these state and action choices are very effective in existing pipelines, they can disconnect the reasoning about actions from the state representation. When reasoning about how to shape clay, humans determine at a high level where and how the clay needs to be deformed. In this work, we propose a simple gridded top-down image of the clay as the state and action space to allow the LLM to reason about the larger geometric regions of the clay without needing to directly reason about global end effector poses.
@@
III-B Prompting
The goal of this work is to explore how we can leverage the world knowledge and shape-level reasoning capabilities of LLMs without requiring any further training or fine-tuning on our specific task. The full process diagram of the iterative prompting scheme is shown in Figure 2, in which we use an action prompt and a termination prompt. The role of the action prompt is to query the LLM with an explanation of the task, state and action space, and the robot embodiment to then output a crafting action sequence of varying length. We take inspiration from past successful approaches using LLMs for robotic tasks for some of the prompting details [kwon2023language]. The components of the action prompt are as follows: (1) an overview explaining the robot embodiment and the high-level details of the crafting task, (2) an explanation of the environmental setup including the clay and the grid coordinate system, (3) an explanation of the grasp action, including that the end-effector pose cannot change during the grasp, (4) an explanation of how to generate a crafting trajectory, including the requirements of (a) describing the current and goal shapes, (b) reasoning about the grasp strategy, (c) describe the predicted effect of the chosen grasp to the clay state, (5) the goal gridded image, (6) the current state gridded image, and (7) the user command to create the goal from the current state. The role of the termination prompt is to query the LLM to determine if the goal shape has been reached. The advantage of using an LLM as opposed to similarity metric thresholding is that we are prompting the LLM to judge the shape similarity more holistically, i.e. if both of the shapes are of the letter T, which is a fundamentally different task from judging point-wise similarity between point clouds. The terimation prompt consists of (1) an overview of the shape comparison task, (2) an explanation of how to generate a decision, with the requirements to (a) describe the similarities and differences between the state and goal, and (b) make a decision if crafting should continue, (3) the current gridded state image, (4) the goal gridded image, and (5) the user command to determine if crafting should stop. For full details of both prompts, please see the Supplemental Materials. In section IV-E we will discuss prompt ablations to evaluate the impactfullness of each component of the prompts on the final system performance.
IV Experiments
@@ IV-A State Representation
IV-B Single-Step Goals
@@
Before we leverage the LLM for longer-horizon shape crafting goals, we first need to evaluate how well an LLM can out-of-the-box select the single action to achieve a single-step goal. While this is a relatively simple task, it proved somewhat difficult for the LLM due to how it is reasoning about the clay moulding task. We found that the LLM prescribes semantically meaningful descriptions for the clay states and goals in order to generate action sequences. This is a very useful attribute for the long horizon shape crafting task, but can interfere with performance for the single-step goals in which the goal shape isn’t anything in particular. For more detailed analysis of the LLM reasoning, please see the Supplemental Materials. We found the LLM was able to successfully reach the goal shape, but could sometimes get confused about rotation. The numerical results for the single-step experiments are shown in Table I, and the qualitative results are visualized in Figure 4.
IV-C Long-Horizon Shape Goals
@@
To further investigate the capabilities of LLMs for crafting, we evaluated the full system performance on a set of long-horizon letter shape goals consisting of the shapes C, I, L, T, X. We selected these letters because they are all able to be created with only deformation actions, whereas letters such as P are not because our current action formulation cannot create holes. To determine the effectiveness of the proposed iterative prompting scheme, we compared the performance with replanning enabled compared to executing the entire LLM-generated trajectory without replanning. Additionally, we compare the iterative LLM-Craft with an iterative variant in which the LLM also controls how much the end-effector is squeezed with options ’min’, ’medium’ and ’max’. Finally, to disambiguate the system errors from the limitations of the coarsely discretized action space, we conducted a set of human baseline experiments in which the human operator must select the two cells to squeeze together. The qualitative results are visualized in Figure 5, and the quantitative results are shown in Table II. Beyond presenting the CD/EMD similarity metrics with respect to the goal, we conduct human evaluation surveys and present the human classification accuracy for the shapes created by each experiment variant. For more details of how the human evaluation surveys were conducted, please see the appendix. We find that the LLM-Craft framework is able to successfully create the simple letter shaped sculptures at a comparable level to our human baseline with the discrete action space. Additionally, LLM-Craft can create smoother shapes when it also controls how much the gripper squeezes the clay. However, it is important to note that the performance of LLM-Craft has a high variance, as sometimes the system selects a bad action early into the trajectory and even with the iterative re-planning framework, the system is unable to recover.
COULD INCLUDE EXPERIMENTS FOR TEXT-ONLY GOALS!
TODO: get the squeeze variant human eval results
TODO: update the project website
| Loop | CD | EMD | Human Eval. |
| No Replan | 0.0056 0.0019 | 0.0082 0.0026 | 0.63 0.15 |
| Iterative | 0.0045 0.0015 | 0.0067 0.0020 | 0.59 0.33 |
| Squeeze | 0.0055 0.0022 | 0.0078 0.0025 | - |
| Text Goal | - | - | - |
| Human | 0.0051 0.0029 | 0.0082 0.0037 | 0.78 0.11 |
| CD | EMD | ||
| 6x6 grid | LLM-Craft | 0.0061 0.0009 | 0.0078 0.0011 |
| Human | 0.0054 0.0007 | 0.0073 0.0004 | |
| 8x8 grid | LLM-Craft | 0.0042 0.0010 | 0.0055 0.0009 |
| Human | 0.0041 0.0007 | 0.0056 0.0005 | |
| 10x10 grid | LLM-Craft | 0.0053 0.0005 | 0.0073 0.0009 |
| Human | 0.0038 0.0005 | 0.0055 0.0007 | |
| 12x12 grid | LLM-Craft | 0.0062 0.0004 | 0.0083 0.0010 |
| Human | 0.0041 0.0008 | 0.0061 0.0005 | |
| 16x16 grid | LLM-Craft | 0.0060 0.0016 | 0.0084 0.0017 |
| Human | 0.0039 0.0010 | 0.0054 0.0011 |
IV-D Grid Ablation
To evaluate if the proposed framework can scale successful with different grid sizes, we deployed the LLM-Craft iterative pipeline on hardware for the ‘X’ letter task with grids of 6x6, 8x8, 10x10, 12x12 and 16x16. We conduct 5 experimental runs for each grid and present the visualization of the shape results in Figure 6 and the quantitative results in Table III. Additionally, we extended the human baseline in which a human selects actions with the same discrete action space to better visualize the shape capabilities of the constrained state/action space. For this ablation experiment, we kept the 3D printed grippers the same across experiments for the sake of time, meaning the varying grid size allowed for finer control of end-effector position, but the same quantity of clay was still being moved with each squeeze. Given the results of these experiments, we believe that the level of detail of shapes created could be more refined with the combination of these finer grids and thinner fingertips.
IV-E Prompt Ablation
@@
To evaluate the impact of different components of the prompts on the system performance, we conducted an ablation study removing different parts. In particular, we chose to remove the following: (1) asking the LLM to predict the effect of the grasp on the clay after selecting an action, (2) asking the LLM to describe the similarities and differences between the state and goal before generating an action sequence, (3) removing the termination prompt altogether, (4) removing the goal text description, i.e. if the goal is an L to not mention L in the prompt explicitly, (5) removing the explanation of grasp types, specifically horizontal, vertical, or diagonal, (6) removing the instructions to choose action step-by-step, and (7) removing the explanations of the gridded state and action space. For the complete prompt and prompt variations for each ablation experiment, please see the Supplemental Materials. The breakdown of earth mover’s distance for the L shape across the ablations is shown in Figure 7. We chose to conduct the ablation experiments with the L shape goal because it is difficult to successfully create the bottom perpendicular line in the L given the discrete action space. Therefore, we believe this goal shape will help illuminate behavior differences given the changes to the prompt. Based on our ablation experiments, it is clear that removing the explanation of the grid from the prompt negatively impacts performance the most. This is an expected result, as the choice of state and action space in LLM-Craft is unique and requires a clear explanation, particularly of the labeling system of the grid in order to select the grid cells for actions. Removing the step-by-step instruction also significantly impacted performance, following the results from [kwon2023language]. We found that the grasp orientation description, goal text description and the termination prompt all had a mid-sized impact on the final performance of LLM-Craft. Finally, the requirements to reason about the state and goal as well as the predicted effect before selecting a grasp action had minimal impact on the average performance of LLM-Craft, but the variance was substantially reduced with their inclusion in the prompt. This makes sense, as requiring the LLM to reason about the differences between the current state and goal as well as describe physically how a grasp will change the state has the potential to reduce the frequency of physically unreasonable actions to be chosen. However, it is important to highlight that while all these components of the prompt reduce the unreasonable actions selected by the LLM, LLM-Craft does sometimes select actions and the explanation of the predicted effect is not fully physically accurate.
IV-F Semantic Goals